OpenStudy (anonymous):

What polynomial has roots of -6, -4, and 1 x^3 - 9x^2 -22x +24 x^3 -x^2 -26x -24 x^3 +x^2 -26x +24 x^3 +9x +14x -24

OpenStudy (anonymous):

put in the polynomial x=1 ,x=-4,x=-6 for that polynomial for for which you get zero is the root. for example put x=1 in x^3 - 9x^2 -22x +24 (1)^3-9(1)^2-22(1)+24=-6 since i am not getting zero so x=1 is not a root. try the same for the above all polynomials.

OpenStudy (anonymous):

(x+6)*(x+4)*(x-1)=0 - this equation can be formed from the given roots. Simplify it to get the answer x^3 +9x^2 +14x -24 = 0

OpenStudy (anonymous):

@sami-21 's method also works well but takes a little bit more time.

OpenStudy (shubhamsrg):

sum of roots = -b/a sum of product in pair = c/a and product = -d/a find these coefficients and hence you have the eqn.. for simplicity, take a=1 ..

OpenStudy (shubhamsrg):

just by checking for -b/a ,,you have your ans straightaway,, hope am explaining this clear enough,,